### v3

# Progressive Binarization with Semi-Structured Pruning for LLMs

Abstract
Large language models (LLMs) have achieved remarkable progress in natural language processing, but their high computational and memory costs hinder deployment on resource-constrained devices. Binarization, which reduces model weights to 1 bit, is a promising solution for efficient inference. However, binarized LLMs still exhibit redundancy that can be further compressed. Semi-structured pruning offers a favorable trade-off between model performance and hardware efficiency, but naively combining it with binarization often leads to severe performance degradation. To address this, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P), a novel post-training compression framework. We propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO) to jointly reduce pruning and binarization error. Additionally, we develop a Coarse-to-Fine Search (CFS) strategy to more effectively select pruning elements. Extensive experiments across multiple LLM families show that PBS2P consistently outperforms state-of-the-art binary post-training quantization methods in both perplexity and downstream accuracy. The code and models will be available at https://github.com/XIANGLONGYAN/PBS2P.
1 Introduction
Figure 1: Perplexity of LLaMA-13B on WikiText2 under different bit-widths across various methods.
Transformer-based large language models (LLMs) vaswani2017attention have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks. This success is largely driven by their massive scale, often comprising billions of parameters. For instance, the OPT (Open Pre-trained Transformer) series zhang2022opt includes models with up to 66 billion parameters, while the LLaMA family touvron2023llama1 features even larger models, such as LLaMA3-70B dubey2024llama3 . Although these models have substantially advanced the state-of-the-art (SOTA) in NLP, their increasingly computational and memory demands present significant challenges for deployment on the resource-constrained devices.
The compression of LLMs can be broadly categorized into several approaches, including weight quantization lin2024awq ; frantar2022gptq , low-rank factorization zhang2023loraprune ; yuan2023asvd , network pruning sun2023simple ; frantar2023sparsegpt , and knowledge distillation zhong2024revisiting ; gu2023knowledge . Among these, binarization represents an extreme form of quantization, reducing model weights to 1 bit and drastically lowering memory consumption. Recent efforts primarily focus on post-training quantization (PTQ) huang2024billm ; li2024arb ; dong2024stbllm , which eliminates the need for backpropagation, thereby accelerating the binarization process and improving deployment efficiency. For example, BiLLM huang2024billm introduces a residual approximation strategy to enhance the performance of 1-bit LLMs, while ARB-LLM li2024arb adopts an alternating refinement approach to better align binarized weights with their full-precision counterparts. STBLLM dong2024stbllm further advances this line of work by compressing LLMs to sub-1-bit precision. Despite the aforementioned advancements, binarized LLMs continue to exhibit redundancy, indicating opportunities for additional compression.
Pruning lecun1989optimal is a promising technique for further reducing redundancy in binarized models. However, conventional structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation often leads to severe performance degradation in LLMs. On the other hand, unstructured pruning dong2024pruner suffers from poor hardware compatibility and inefficient storage. Semi-structured pruning frantar2023sparsegpt ; sun2023simple ; dong2024stbllm has emerged as an effective compromise, significantly reducing redundancy while maintaining a balance between accuracy and hardware efficiency. Nonetheless, directly combining binarization with semi-structured pruning often leads to noticeable performance degradation, making it hard to maintain model accuracy. Another challenge is how to select pruning elements effectively to improve pruning efficiency while preserving accuracy.
To address these challenges, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLMs, which achieves substantial model compression while maintaining strong performance (see Figure 1). We first propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively prunes a subset of elements at each step while jointly optimizing the binarized parameters. This strategy effectively reduces the cumulative error from pruning and binarization. To further enhance pruning efficiency, we develop a Coarse-to-Fine Search (CFS) algorithm for more accurate pruning element selection. In the coarse stage, pruning ratios are assigned to each layer based on its importance. In the fine stage, a Hessian-based metric is used to identify the most redundant elements, guided by the layer-specific pruning ratios. We further show that the Hessian-based metric minimizes the theoretical error increase under a second-order approximation. This makes it a provably optimal strategy for pruning.
Extensive experiments demonstrate that PBS2P consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing binary PTQ methods on a wide range of evaluation benchmarks. As shown in Figure 1, on the WikiText-2 merity2016pointer dataset, PBS2P attains a perplexity of 6.20 on LLaMA-13B touvron2023llama1 with an average bit-width of just 0.8 bits, compared to 5.47 for the full-precision model. These results further highlight the effectiveness of PBS2P in significantly narrowing the performance gap between binarized and full-precision models.
Our key contributions can be summarized as follows:
  • We propose a novel framework, PBS2P, which integrates binarization and semi-structured pruning seamlessly for effective LLM compression.
  • We propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively prunes the model while jointly optimizing binarization parameters, effectively reducing the combined error and preserving performance.
  • We propose Coarse-to-Fine Search (CFS) for selecting pruning elements, which enables effective identification of redundant parameters while better preserving model performance.
  • Extensive experiments show that PBS2P consistently outperforms SOTA binary PTQ methods and significantly narrows the gap to full-precision models across diverse benchmarks.
2 Related works
2.1 LLM Quantization
Quantization compresses full-precision parameters into lower-bit representations, reducing both computation and storage demands. Current quantization methods for LLMs are mainly divided into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT liu2023llm ; chen2024db ; du2024bitdistiller integrates quantization during the training phase to enhance low-bit weight representations. However, due to the enormous parameter number, retraining becomes excessively expensive and inefficient for LLMs. PTQ, as it directly applies quantization to the model weights without retraining, making it faster and less resource-demanding. Recent methods, like ZeroQuant yao2022zeroquant and BRECQ li2021brecq , improve quantization accuracy by incorporating custom quantization blocks and group labels. While GPTQ frantar2022gptq and QuIP chee2024quip use second-order error compensation to reduce quantization errors.
2.2 Network Binarization
Binarization, as the most extreme form of quantization, reduces model parameters to a single bit (±1). Prominent methods, like Binary Weight Network (BWN) rastegari2016xnor , XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi , focus on binarizing the weights, with XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi also binarizing activations. In the context of LLM binarization, BitNet wang2023bitnet , OneBit (xu2024onebit, ), and BinaryMoS jo2024mixture adopt the QAT framework, while BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLMdong2024stbllm use PTQ combined with residual approximation. Our work focuses on the binary PTQ method, achieving substantial improvements over existing SOTA binary PTQ methods.
2.3 Model Pruning
Pruning lecun1989optimal is a common technique for compressing neural networks by removing less important parameters, resulting in smaller and more efficient sparse models. In LLMs, pruning methods are typically categorized as structured, unstructured, or semi-structured. Structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation removes entire components to enhance efficiency but often causes performance drops, requiring retraining. Unstructured pruning dong2024pruner eliminates individual weights based on importance, preserving performance even at high sparsity, though the irregular patterns are hardware-unfriendly. Semi-structured pruning offers a good trade-off, enforcing regular patterns like $N$:$M$ sparsity for hardware efficiency, as used in SparseGPT frantar2023sparsegpt , Wanda sun2023simple , and STBLLM dong2024stbllm . Our method adopts semi-structured $N$:$M$ sparsity to retain performance while ensuring hardware compatibility.
Figure 2: Overview of our PBS2P framework. Coarse-Stage Search: using the layer importance to assign pruning ratios to each layer. Fine-Stage Search: searching the elements to be pruned based on the Hessian-based metrics. Stepwise semi-structured Pruning with Binarization Optimization: stepwise pruning with optimization of binarized parameters.
3 Methodology
3.1 Preliminary
Binarization.    First, let’s briefly review the process of standard binarization rastegari2016xnor . Binarization begins by performing a row-wise redistribution on the full-precision matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ , resulting in a matrix $\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}$ with a mean of 0:
$\widetilde{\mathbf{W}}=\mathbf{W}-\mu,\quad\text{where}\,\mu=\frac{1}{m}\sum_{j=1}^{m}\mathbf{W}_{.j}\,.$ (1)
The binarization objective is to minimize the reconstruction error: $\arg\min_{\alpha,\mathbf{B}}\|\widetilde{\mathbf{W}}-\alpha\mathbf{B}\|_{F}^{2}$, where $\alpha\in\mathbb{R}^{n}$ is a row-wise scaling factor and $\mathbf{B}\in\{+1,-1\}^{n\times m}$ is the binary matrix. The optimal solutions for $\alpha$ and $\mathbf{B}$ are given by $\alpha=\frac{1}{m}\sum_{j=1}^{m}|\widetilde{\mathbf{W}}_{.j}|$ and $\mathbf{B}=\operatorname{sign}(\widetilde{\mathbf{W}})$, respectively huang2024billm .
N:M Sparsity.    $N$:$M$ sparsity is a semi-structured pruning scheme that balances model compression and hardware efficiency. Unlike unstructured pruning, which removes arbitrary weights and results in irregular sparsity patterns that are hard to accelerate, $N$:$M$ sparsity imposes a structured constraint by retaining exactly $N$ nonzero elements out of every $M$ consecutive weights frantar2023sparsegpt ; sun2023simple ; dong2024stbllm . A typical configuration, such as 2:4 sparsity, keeps only two weights out of every four, significantly reducing parameter count while preserving computational efficiency. This regular pattern allows for efficient execution on NVIDIA Ampere GPUs nvidia2020 , taking advantage of specialized hardware support. In our work, we adopt the $N$:$M$ sparsity scheme as a semi-structured pruning strategy to further eliminate redundancy within the binarized model. This combination enables more aggressive compression while maintaining hardware compatibility and inference efficiency.
3.2 Stepwise Semi-Structured Pruning with Binarization Optimization
We define the total error $\mathcal{L}$ for binarization and semi-structured pruning as follows:
$\mathcal{L}=||\mathbf{W}\mathbf{X}-(\mathbf{M}_{\text{p}}\odot\mathbf{\widehat{W}})\mathbf{X}||_{F}^{2},\quad\text{where}\,\widehat{\mathbf{W}}=\alpha\mathbf{B}+\mu\,.$ (2)
where $\mathbf{X}$ denotes the calibration data, $\mathbf{M}_{\text{p}}$ denotes the semi-structured pruning mask. Optimizing both the pruning mask $\mathbf{M}_{\text{p}}$ and the binarized parameters $\mathbf{\widehat{W}}$ simultaneously is an NP-hard problem blumensath2008iterative . Therefore, we adopt a greedy-style approximation. At each step, we prune a small portion of elements. Then, we update the binarized parameters to fit the current pruning ratio. This step-by-step strategy reduces the optimization difficulty. It distributes the optimization burden across multiple pruning stages. The selection of the semi-strucutred pruning mask $\mathbf{M}_{\text{p}}$ will be elaborated in Section 3.3.
Figure 3: Illustration of the $N$:$M$ sparsity pruning mask group, where the original pruning mask $\mathbf{M}_{\text{p}}$ is divided into $M$-$N$ sub masks.
N:M Sparsity Pruning Mask Group.
To enable progressive pruning, we decompose the original pruning mask $\mathbf{M}_{\text{p}}$ into a series of sub-masks, each representing a step of pruning. Under the $N$:$M$ sparsity constraint, exactly $N$ out of every $M$ elements are retained. Within each $M$-element group, the $M{-}N$ pruned elements are removed one by one across successive steps. Formally, the original pruning mask $\mathbf{M}_{\text{p}}$ can be expressed as the union of $M{-}N$ sub-masks:
$\mathbf{M}_{\text{p}}=\bigcup_{i=1}^{M-N}\mathbf{M}_{\text{p}}^{i}.$ (3)
Figure 3 illustrates this process. This decomposition supports progressive pruning, facilitating tighter integration with binarization in subsequent stages.
Stepwise Pruning with Binarization Optimization (SPBO).    After obtaining the group of pruning masks, we proceed with the stepwise pruning and binarization optimization process. We perform a total of $M{-}N$ pruning steps to achieve $N$:$M$ sparsity. After each step, we update the binarization parameters $\alpha$ and $\mu$ to better adapt to the current pruning state. At the $k$-th pruning step, we obtain the updated binary matrix $\mathbf{B}_{k}$ and full-precision matrix $\mathbf{W}_{k}$ by applying the $k$-th pruning sub-mask $\mathbf{M}_{\text{p}}^{k}$ to the previously pruned matrices $\mathbf{B}_{k-1}$ and $\mathbf{W}_{k-1}$:
$\mathbf{B}_{k}=\mathbf{B}_{k-1}\odot\mathbf{M}_{\text{p}}^{k},\quad\mathbf{W}_{k}=\mathbf{W}_{k-1}\odot\mathbf{M}_{\text{p}}^{k}.$ (4)
After the $k$-th pruning step, the binarization parameters $\alpha$ and $\mu$ are no longer optimal under the updated sparsity pattern. To minimize the error $\mathcal{L}$ in the current pruning state, these parameters must be re-estimated. Inspired by ARB-LLM li2024arb , a straightforward approach is to compute the partial derivatives of $\mathcal{L}$ with respect to $\alpha$ and $\mu$, and solve for the stationary points by setting them to zero. Specifically, we first update $\mu$ by setting $\partial\mathcal{L}/\partial\mu=0$ during the $k$-th pruning step:
$\partial\mathcal{L}/\partial\mu=0\quad\Rightarrow\quad\mu=\frac{\mathbf{1}^{\top}\mathbf{S}(\mathbf{W}_{k}-\alpha\mathbf{B}_{k})^{\top}}{\mathbf{1}^{\top}\mathbf{S}\mathbf{1}},$ (5)
where $\mathbf{S}=\sum\nolimits_{b}\mathbf{X}_{b}^{\top}\mathbf{X}_{b}$ is a precomputed matrix derived from the calibration data $\mathbf{X}$ for notational convenience. Then, we adjust $\alpha$ by setting $\partial\mathcal{L}/\partial\alpha=0$:
$\partial\mathcal{L}/\partial\alpha=0\quad\Rightarrow\quad\alpha=\frac{\operatorname{diag}(\mathbf{B}_{k}\mathbf{S}(\mathbf{W}_{k}-\mu)^{\top})}{\operatorname{diag}(\mathbf{B}_{k}\mathbf{S}\mathbf{B}_{k}^{\top})}.$ (6)
During the $k$-th pruning round, we can alternately update $\mu$ and $\alpha$ for $\tau$ steps to further reduce $\mathcal{L}$. We perform total $M$-$N$ steps of pruning until all pruning elements are removed. The detailed derivation of optimizing $\mu$ and $\alpha$ can be found in the supplementary material.
3.3 Coarse-to-Fine Search
Selecting effective pruning elements is a key challenge in model compression. To address this, we adopt a two-stage strategy. In the coarse stage, we assign pruning ratios to each layer based on its layer importance. In the fine stage, we determine the specific elements to prune within each layer according to the assigned ratios, allowing for more accurate selection of pruning candidates.
Coarse-Stage: Pruning Ratio Allocation.    To estimate the importance of each layer, we adopt a simple criterion that quantifies how much a layer influences the input-output relationship. Layers that induce greater changes between input and output are considered more important, whereas those with limited impact are deemed less important. Specifically, for a given weight layer $\mathcal{W}$, let $\mathcal{X}_{\mathcal{W}}$ and $\mathcal{Y}_{\mathcal{W}}$ denote the input and output, respectively. The importance of the layer is then evaluated as follows:
$\text{Importance}(\mathcal{W})=\frac{1}{\text{similarity}(\mathcal{X_{\mathcal{W}},Y_{\mathcal{W}}})},\quad\text{where}\,\,\,\text{similarity}(\mathcal{X_{\mathcal{W}},Y_{\mathcal{W}}})=\frac{\mathcal{X}_{\mathcal{W}}\cdot\mathcal{Y}_{\mathcal{W}}}{\|\mathcal{X}_{\mathcal{W}}\|_{2}\|\mathcal{Y}_{\mathcal{W}}\|_{2}}.$ (7)
Inspired by dumitru2024change , we employ cosine similarity to assess the alignment between input and output representations. Based on the computed importance, we rank all layers, with $k_{i}$ denoting the rank of the $i$-th layer. The pruning parameter $N_{i}$ for each layer is determined as follows:
$N_{i}=\left\lfloor N_{\text{high}}-\left(N_{\text{high}}-N_{\text{low}}\right)\cdot\frac{k_{i}-1}{L-1}+\frac{1}{2}\right\rfloor,$ (8)
where $L$ represents the total number of layers in the model. $N_{\text{high}}$ is the maximum pruning parameter, and $N_{\text{low}}$ is the minimum pruning parameter. Both are predefined constants. For example, suppose the $m$-th layer is the most important with $k_{m}=1$. It will be assigned $N_{\text{high}}$, retaining more elements. In contrast, if the $n$-th layer is the least important with $k_{n}=L$, it will be assigned $N_{\text{low}}$, preserving the fewest elements. We define the average pruning parameter $N_{\text{average}}$ as the average of the maximum and minimum pruning parameters, such that $N_{\text{average}}=(N_{\text{high}}+N_{\text{low}})/2$.
Fine-Stage: Selecting Pruning Elements.    Once the pruning ratio $N/M$ is assigned in the coarse-stage, we proceed to search for the specific pruning elements within each layer. We revisit the impact of pruning on model error. Inspired by OBS hassibi1993optimal , we derive the relationship between pruning and error increase under a second-order approximation, and present the corresponding formulation in Theorem 3.1. The full proof is provided in the supplementary material. Since our objective is to minimize the increase in error caused by pruning, we directly adopt the following score to guide element selection: $s_{q}=\frac{w_{q}^{2}}{\left[\mathbf{H}^{-1}\right]_{qq}^{2}}$. This criterion is theoretically guaranteed to minimize the error increase introduced by pruning under a second-order approximation.Based on the given $N$:$M$ ratio and the corresponding $s_{q}$ values, we select the top $N$ elements with the largest $s_{i}$ values from every $M$ consecutive elements to retain. The other $M-N$ elements are selected for pruning.
Theorem 3.1.
After pruning an element $w_{q}$, the pruning-induced error increase $\delta\mathcal{L}$ can be approximated by the following expression: $\delta\mathcal{L}=\frac{1}{2}\frac{w_{q}^{2}}{[\mathbf{H}^{-1}]_{qq}},$ (9) where $[\mathbf{H}^{-1}]_{qq}$ denotes the $q$-th diagonal entry of the inverse Hessian matrix.
Algorithm 1 Pseudocode of PBS2P
1:Inputs: Model $\mathcal{M}$, Calibration Data $\mathcal{X}$, Target Pruning Ratio $\mathcal{TR}$, Optimization Steps $\tau$
2:Output: Binarized and Pruned Model $\mathcal{M}^{\prime}$
3:procedure PBS2P($\mathcal{M},\mathcal{X},\mathcal{TR},\tau$)
4:     $R\leftarrow$ Coarse_Stage_Search$(\mathcal{M},\mathcal{X},\mathcal{TR})$triangleright Allocate layer-wise pruning ratios
5:     for each layer $l$ in model $\mathcal{M}$ do
6:         $\mathcal{W}_{l}\leftarrow$ Extract_Weights$(\mathcal{M},l)$triangleright Obtain weights of layer $l$
7:         $r_{l}\leftarrow R[l]$triangleright Get pruning ratio for layer $l$
8:         $\mathbf{M}_{\text{p}}\leftarrow$ Fine_Stage_Search$(\mathcal{W}_{l},\mathcal{X},r_{l})$triangleright Generate pruning mask
9:         $\mathcal{W}_{l}^{\prime}\leftarrow$ SPBO$(\mathcal{W}_{l},\mathcal{X},\mathbf{M}_{\text{p}},\tau)$triangleright Stepwise pruning and binarization optimization
10:         $\mathcal{M}^{\prime}\leftarrow$ Update_Layer$(\mathcal{M},l,\mathcal{W}_{l}^{\prime})$triangleright Write updated weights back to model
11:     end for
12:     return $\mathcal{M}^{\prime}$
13:end procedure
3.4 PBS2P Pipeline
PBS2P Workflow.    As illustrated in Figure 2, the proposed PBS2P framework is composed of three key components: Coarse-stage Search, Fine-stage Search, and Stepwise Semi-structured Pruning with Binarization Optimization. The workflow of PBS2P is outlined in Algorithm 1. Given a full-precision model $\mathcal{M}$, calibration data $\mathcal{X}$, and a target average pruning ratio $\mathcal{TR}$, the algorithm first performs a coarse-stage search to determine layer-wise pruning ratios based on importance. With the ratios determined, we proceed layer by layer to apply pruning and binarization. For each layer $l$, we extract its weight matrix $\mathcal{W}_{l}$ and conduct a fine-stage search to generate a semi-structured pruning mask $\mathbf{M}_{\text{p}}$. The SPBO procedure is then invoked to progressively prune the weights while optimizing the binarization parameters. This joint optimization reduces the total error introduced by pruning and binarization. The updated weights $\mathcal{W}_{l}^{\prime}$ are written back to the model, and the process repeats for all layers. As pruning and binarization are applied layer-wise, our method has low memory overhead.
Average Bits.    Following BiLLM huang2024billm and STBLLM dong2024stbllm , we similarly divide weights into salient and non-salient groups, consistent with their design. For salient weights, we adopt a residual binarization strategy, allocating two bits to better preserve their expressiveness, as done in both prior works. For non-salient weights, we apply group-wise quantization. In line with STBLLM dong2024stbllm , we further divide the non-salient weights into three groups using two split points, maintaining the same quantization scheme. The weight parameters and additional hardware overhead are as follows:
$\left\{\begin{array}[]{lr}N_{\text{param}}=\left[2\times r_{\text{salient}}+(1-r_{\text{salient}})\right]\times\frac{N}{M},\\ N_{\text{storing}}=2+\dfrac{1}{b_{\text{size}}},\\ \end{array}\right.$ (10)
where $r_{\text{salient}}$ represents the proportion of salient weights, $\frac{N}{M}$ denotes the predefined avrage pruning ratio for the entire model, and $b_{\text{size}}$ indicates the block size used in OBC frantar2022optimal compensation, with 2 bits reserved to mark the division between salient and non-salient weights. Our parameter settings and $\frac{N}{M}$ configuration are identical to our main comparison method STBLLM dong2024stbllm . Specifically, we adopt the same settings of 4:8, 5:8, and 6:8 sparsity as used in STBLLM dong2024stbllm . These correspond to average bit-widths of 0.55, 0.70, and 0.80, respectively.
4 Experiments
4.1 Settings
All experiments are performed using PyTorch paszke2019pytorch and Huggingface paszke1912imperative on a single NVIDIA A800-80GB GPU. Following the work of frantar2022gptq , huang2024billm , and li2024arb , we use 128 samples from the C4 (raffel2020exploring, ) dataset for calibration. Since PBS2P is an efficient PTQ framework, it eliminates fine-tuning, enabling completion through a single process combining binarization and pruning.
Models and Datasets.    We conduct extensive experiments on the LLaMA touvron2023llama1 , LLaMA-2 touvron2023llama2 , and LLaMA-3 dubey2024llama3 families and the OPT family (zhang2022opt, ). To evaluate the effectiveness of PBS2P, we measure the perplexity of LLM’s outputs on WikiText2 (merity2016pointer, ), PTB (marcus1994penn, ), and C4 (raffel2020exploring, ). Moreover, we also evaluate the accuracy for 7 zero-shot QA datasets: ARC-c (clark2018think, ), ARC-e (clark2018think, ), BoolQ (clark2019boolq, ), Hellaswag (zellers2019hellaswag, ), OBQA (mihaylov2018can, ), RTE (chakrabarty2021figurative, ), and Winogrande (sakaguchi2019adversarial, ).
Baselines.    We mainly compare our PBS2P with STBLLM dong2024stbllm , a structural binary PTQ framework designed for compressing LLMs to precisions lower than 1-bit. We compare the results of PBS2P with STBLLM under the same $N$:$M$ settings (e.g., 4:8, 5:8, 6:8). Previous low-bit methods like ARB-LLM li2024arb , BiLLM huang2024billm , GTPQ (frantar2022gptq, ) and vanilla RTN are also selected for comparison.
Table 1: Perplexity comparison of RTN, GPTQ frantar2022gptq , BiLLM huang2024billm , ARB-LLM li2024arb , STBLLM dong2024stbllm , and PBS2P on the LLaMA families. The evaluation results demonstrate the perplexity performance on the Wikitext2 (merity2016pointer, ) dataset across various model sizes.
Settings LLaMA-1 LLaMA-2 LLaMA-3 Method #Block W-Bits 7B 13B 30B 65B 7B 13B 8B FP16 - 16 5.68 5.09 4.10 3.53 5.47 4.88 6.14 RTN - 3 25.54 11.40 14.89 10.59 542.80 10.68 2194.98 GPTQ 128 3 8.63 5.67 4.87 4.17 6.44 5.46 18.68 RTN - 2 106767.34 57409.93 26704.36 19832.87 17788.94 51145.61 1335816.13 GPTQ 128 2 129.20 20.46 15.29 8.66 52.22 23.63 1480.43 RTN - 1 168388.00 1412020.25 14681.76 65253.24 157058.34 47902.32 1353698.38 GPTQ 128 1 164471.78 131505.41 10339.15 20986.16 59758.69 22926.54 1121260.50 BiLLM 128 1.11 49.79 14.58 9.90 8.37 32.31 21.35 55.80 ARB-LLM 128 1.11 14.03 10.18 7.75 6.56 16.44 11.85 27.42 STBLLM 128 0.80 15.03 9.66 7.56 6.43 13.06 11.67 33.44 STBLLM 128 0.70 19.48 11.33 9.19 7.91 18.74 13.26 49.12 STBLLM 128 0.55 31.72 17.22 13.43 11.07 27.93 20.57 253.76 PBS2P 128 0.80 7.36 6.20 5.21 4.60 7.17 6.27 10.75 PBS2P 128 0.70 8.09 6.85 5.78 5.09 8.00 6.89 12.29 PBS2P 128 0.55 10.78 9.24 7.19 6.39 10.64 8.68 17.45
4.2 Main Results
We perform a comprehensive comparison of different LLM families (like LLaMA touvron2023llama1 ; dubey2024llama3 ; touvron2023llama2 and OPT zhang2022opt ) with various model sizes. To keep fairness, We follow STBLLM dong2024stbllm to report the average bit-width of all methods, where our methods have the same bit-width as STBLLM.
Results on LLaMA Family.   As shown in Table 1, the models using RTN and GPTQ methods find it hard to maintain model performance at 1-bit precision. BiLLM achieves a satisfactory perplexity of 1.11 bits but performs worse than ARB-LLM at the same bit-width. At sub-1-bit precision, PBS2P surpasses STBLLM and significantly reduces perplexity at the same bit-width across model sizes from 7B to 65B. For instance, PBS2P achieves a substantial improvement over STB-LLM on LLaMA-1-7B, with perplexity dropping from 31.72 to 10.78, a reduction of approximately 66.0%, in the extreme case of 4:8 structured binarization, where half of the parameters are pruned.
Table 2: Perplexity on the OPT family using Wiki2.
Method Block W-Bit 1.3B 2.7B 30B FP16 - 16 14.62 12.47 9.56 RTN - 2 11272.65 9505.76 1165864.25 GPTQ 128 2 121.64 59.53 13.04 RTN - 1 17165.72 36516.69 6485.99 GPTQ 128 1 8719.58 11700.13 14083.15 BiLLM 128 1.11 69.05 48.61 13.86 ARB-LLM 128 1.11 26.63 19.84 11.12 STBLLM 128 0.80 29.84 17.02 12.80 STBLLM 128 0.70 33.01 20.82 14.38 STBLLM 128 0.55 45.11 30.34 18.80 PBS2P 128 0.80 23.72 18.32 10.50 PBS2P 128 0.70 27.10 20.17 10.82 PBS2P 128 0.55 44.53 27.42 11.87
Furthermore, PBS2P, with a precision of 0.8 bits, outperforms both RTN at 3 bits, GPTQ at 2 bits, BiLLM and ARB-LLM at 1.11 bits, and STBLLM at 0.8 bits in terms of perplexity across all model sizes. Those comparisons show that our PBS2P achieves a better trade-off between bit precision and performance. It is worth noting that PBS2P outperforms GPTQ at 3 bits on LLaMA 1-7B and LLaMA 3-8B. We extend perplexity evaluation to the PTB and C4 datasets. Figure 4 shows the performance of the LLaMA-7B, LLaMA-13B, and LLaMA-2-7B models. PBS2P continues to achieve a leading edge in performance while operating at a relatively lower bit-width compared to other methods.
Results on OPT Family.    We extend our experiments to the OPT family (1.3B to 30B) under sub-1-bit PTQ settings, similar to the setup for LLaMA family. As shown in Table 2, PBS2P continues to outperform STBLLM across most of the models and $N$:$M$ structured binarization configurations. More results are provided in the supplementary material.
4.3 Zero-Shot Results
To provide a more thorough evaluation of binary LLMs, we extend our experiments to 7 zero-shot datasets and test on models from the LLaMA family: LLaMA-1-13B, LLaMA-2-13B, and LLaMA-1-30B. Each model is evaluated across various compression methods, including full-precision, STBLLM (6:8), STBLLM (4:8), PBS2P (6:8), and PBS2P (4:8). As shown in Table 3, models compressed with PBS2P significantly outperform those compressed with STBLLM in terms of average accuracy. Such comparisons demonstrate that PBS2P provides a more effective solution for compressing LLMs to less than 1-bit, offering improved trade-offs between accuracy and efficiency. This highlights its strong potential for enabling high-performance, low-bit inference in real-world applications.
Table 3: Accuracies (%) for 7 zero-shot tasks from semi-structured binarized LLaMA-1-13B, LLaMA-2-13B, and LLaMA-1-30B with STBLLM and PBS2P.
Models Method W-Bits Winogrande OBQA Hellaswag Boolq ARC-e ARC-c RTE Average LLaMA-1-13B FP16 16 72.69 33.20 59.91 77.89 77.40 46.42 70.40 63.80 STBLLM 0.80 65.98 36.20 63.67 65.38 68.86 34.04 56.68 55.83 STBLLM 0.55 63.06 34.80 52.65 62.48 56.90 28.33 52.71 50.13 PBS2P 0.80 72.77 31.00 54.80 74.71 74.37 42.32 68.23 59.74 PBS2P 0.55 69.30 26.80 46.83 71.56 65.70 32.68 55.96 52.69 LLaMA-2-13B FP16 16 72.22 35.20 60.03 80.55 79.42 48.38 65.34 65.00 STBLLM 0.80 63.93 37.00 57.76 71.53 60.56 31.99 54.15 53.85 STBLLM 0.55 55.88 29.40 44.03 64.31 48.86 26.54 52.71 45.96 PBS2P 0.80 72.45 31.00 54.43 80.61 74.28 42.75 59.21 59.24 PBS2P 0.55 69.85 27.00 47.75 75.50 69.19 35.58 62.82 55.38 LLaMA-1-30B FP16 16 75.77 36.00 63.37 82.69 80.30 52.90 67.15 67.40 STBLLM 0.80 71.59 41.00 69.85 77.37 71.55 41.30 48.01 60.10 STBLLM 0.55 64.01 34.60 56.46 63.06 60.86 31.48 51.99 51.78 PBS2P 0.80 75.93 35.00 59.45 82.14 79.29 47.95 63.18 63.28 PBS2P 0.55 72.14 31.20 53.00 79.76 73.99 41.13 69.31 60.08
Figure 4: BiLLM huang2024billm , ARB-LLM li2024arb , and PBS2P performed on the PTB (marcus1994penn, )and c4 raffel2020exploring datasets, mainly on LLaMA-7B, LLaMA-13B, and LLaMA2-7B, and we found that PBS2P at 0.8 bits outperforms other methods.
4.4 Ablation Study
Ablation for SPBO Strategy.    To validate the effectiveness of our SPBO strategy, we provide the performance of PBS2P with and without its application. As shown in Table 4a, the performance of SPBO surpasses that of the vanilla pruning followed by binarization approach. In the vanilla method, pruning is done in a single step, where all elements are pruned simultaneously, and binarization is applied to the remaining elements afterward. This approach makes it difficult to jointly minimize the combined pruning and binarization errors. In contrast, our SPBO strategy progressively couples these two processes, resulting in a clear performance improvement.
Ablation for Metric in Fine-Stage Search.    The Table 4b presents the performance of different pruning metrics in the fine-stage search. We compare several metrics, including random selection, magnitude-based selection, Wanda, SI, and our Hessian-based metric. Our metric significantly outperforms both random selection and magnitude-based pruning, achieving superior performance compared to Wanda and SI, thereby highlighting the effectiveness of our approach.
Ablation for Metric in Coarse-Stage Search.    Table 4c shows the performance of the coarse-stage search under three conditions: without the coarse-stage search, using the relative importance (RI) metric from STBLLM, and applying our layer importance (LI) score. The results reveal that using the RI metric causes a performance drop compared to the baseline without coarse-stage search. In contrast, our proposed layer importance (LI) score achieves substantial improvements, validating both the necessity of the coarse-stage search and the superiority of our metric.
Ablation for Pruning Type.    In Table 4d, we present the impact of different pruning types on the results, comparing structured pruning, unstructured pruning, and semi-structured pruning under the same pruning ratio (50%). We adopt column pruning for the structured case, elementwise magnitude-based pruning for the unstructured case, and 4:8 sparsity for the semi-structured case. Semi-structured pruning strikes an optimal balance: it maintains hardware efficiency while achieving performance close to unstructured pruning and significantly better than structured pruning.
Ablation for Group Size.    Table 4e presents the results of our ablation study on the group size configuration. It indicates that a smaller group size, meaning finer-grained grouping, leads to better performance. However, this also comes with increased computational and storage costs. To strike a balance between performance and resource efficiency, we select a group size of 128.
Ablation for Calibration Data Size.    Similar to other PTQ-based methods, our PBS2P framework requires only a small calibration set. Table 4f shows the effect of varying calibration size. With the C4 dataset, reducing the sample size to 32 or 64 degrades performance, while increasing to 256 yields minimal additional benefit over 128. These results confirm that PBS2P remains robust with just 128 calibration samples.
Table 4: Ablation study on LLaMA-7B, where all PBS2P is applied an $N$:$M$ sparsity of 4:8. Results are measured by perplexity on the Wikitext2 and C4 datasets. Our results are highlighted in bold.
SPBO Strategy Wikitext2$\downarrow$ C4$\downarrow$
14.43 16.76
10.78 13.06
(a) Ablation for SPBO Strategy.
Metric Random Magnitude Wanda SI Ours
Wikitext2$\downarrow$ 7,779.28 38.90 10.89 196.61 10.78
C4$\downarrow$ 6,797.09 24.47 13.16 148.85 13.06
(b) Ablation for Metric in Fine-Stage Search
Coarse-Stage Search Metric Wikitext2$\downarrow$ C4$\downarrow$
11.55 13.94
RI 14.43 16.76
LI 10.78 13.06
(c) Ablation for Metric in Coarse-Stage Search
Prune Type Hardware-friendly Wikitext2$\downarrow$ C4$\downarrow$
Structured 621.16 400.50
Unstructured 8.54 10.95
Semi-Structured 10.78 13.06
(d) Ablation for Pruning Type
Group Size 32 64 128 256 512
Wikitext2$\downarrow$ 9.74 10.24 10.78 11.20 11.91
C4$\downarrow$ 11.75 12.50 13.06 13.68 14.51
(e) Ablation for Group Size
Calibration Data Size 32 64 128 256
Wikitext2$\downarrow$ 11.62 11.20 10.78 10.77
C4$\downarrow$ 13.58 13.44 13.06 13.12
(f) Ablation for Calibration Data Size
4.5 Time Analysis
As a binary PTQ framework, PBS2P eliminates the need for fine-tuning and retraining. It introduces coarse-to-fine search and stepwise pruning with binarization optimization, which moderately increases the overall computation time. As shown in Table 5, BiLLM huang2024billm completes quantization on LLaMA-7B in 45 minutes, while ARB-LLM li2024arb takes 76 minutes.
Table 5: Time comparison of different PTQ methods on LLaMA-7B.
Method Coarse stage SPBO Total
BiLLM 45 min
ARB-LLM 76 min
PBS2P 2 min 109 min 111 min
PPBS2P finishes in 111 minutes, with only 2 minutes spent on the coarse-stage search and 109 minutes on the fine-stage search with SPBO. This additional overhead is mainly due to iterative parameter updates in the fine stage. This overhead is acceptable since all computations are carried out offline, imposing no additional latency during inference. With this design, PBS2P achieves superior accuracy and highly efficient, practical deployment.
5 Discussion
Limitations.  While PBS2P demonstrates strong performance in both accuracy and model compression, it still faces several limitations. First, we have not yet implemented a custom 1-bit kernel that supports the $N$:$M$ sparsity pattern, which limits our ability to accurately evaluate the end-to-end speedup on real hardware platforms. Second, similar to prior binarization methods such as BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLM dong2024stbllm , our approach introduces additional bits to differentiate between salient and non-salient weights. Although this selective binarization strategy helps preserve model performance, it also incurs extra storage overhead, which partially offsets the compression and efficiency gains brought by binarization.
Societal impacts.  Our proposed PBS2P method enables the efficient deployment of large language models on resource-constrained devices, promoting broader access to AI technologies in real-world applications such as mobile assistants, edge computing, and low-power systems. At present, we foresee no negative societal impacts arising from this work.
6 Conclusion
In this work, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLM compression. Central to our approach is SPBO, a stepwise semi-structured pruning strategy with binarization optimization, which prunes a subset of elements at each step while jointly optimizing the binarized parameters. SPBO effectively mitigates the combined error from pruning and binarization. Additionally, we introduce a Coarse-to-Fine Search (CFS) strategy to enhance pruning element selection, further improving compression efficiency. Extensive experiments demonstrate that PBS2P outperforms SOTA binary PTQ methods, delivering superior accuracy across various LLM families and evaluation metrics. Our method provides an effective solution for deploying LLMs on resource-constrained devices while maintaining high performance, and offers a novel perspective on combining compression techniques for extreme LLM compression.

### v4

@@
 Large language models (LLMs) have achieved remarkable progress in natural language processing, but their high computational and memory costs hinder deployment on resource-constrained devices.
-Binarization, which reduces model weights to 1 bit, is a promising solution for efficient inference.
-However, binarized LLMs still exhibit redundancy that can be further compressed.
-Semi-structured pruning offers a favorable trade-off between model performance and hardware efficiency, but naively combining it with binarization often leads to severe performance degradation.
-To address this, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P), a novel post-training compression framework.
-We propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO) to jointly reduce pruning and binarization error.
-Additionally, we develop a Coarse-to-Fine Search (CFS) strategy to more effectively select pruning elements.
-Extensive experiments across multiple LLM families show that PBS2P consistently outperforms state-of-the-art binary post-training quantization methods in both perplexity and downstream accuracy.
+Binarization represents the most extreme form of quantization, yet binarized models still contain redundancy that can be further removed.
+Pruning provides a natural way to eliminate such redundancy, but naïve combination with binarization often results in severe performance degradation.
+In this paper, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P), a novel post-training framework that seamlessly integrates binarization and semi-structured pruning.
+We first propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively introduces sparsity while optimizing binarization parameters to jointly reduce pruning and quantization error, yielding more stable and accurate compression.
+Additionally, we propose a Coarse-to-Fine Search (CFS) that first allocates pruning ratios and then refines element selection, further enhancing overall performance.
+Extensive experiments across multiple LLM families show that PBS2P consistently outperforms state-of-the-art (SOTA) binary post-training quantization methods in both perplexity and downstream accuracy.
 The code and models will be available at https://github.com/XIANGLONGYAN/PBS2P.
@@
 Figure 1: Perplexity of LLaMA-13B on WikiText2 under different bit-widths across various methods.
-Transformer-based large language models (LLMs) vaswani2017attention have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks.
-This success is largely driven by their massive scale, often comprising billions of parameters.
-For instance, the OPT (Open Pre-trained Transformer) series zhang2022opt includes models with up to 66 billion parameters, while the LLaMA family touvron2023llama1 features even larger models, such as LLaMA3-70B dubey2024llama3 .
-Although these models have substantially advanced the state-of-the-art (SOTA) in NLP, their increasingly computational and memory demands present significant challenges for deployment on the resource-constrained devices.
-The compression of LLMs can be broadly categorized into several approaches, including weight quantization lin2024awq ; frantar2022gptq , low-rank factorization zhang2023loraprune ; yuan2023asvd , network pruning sun2023simple ; frantar2023sparsegpt , and knowledge distillation zhong2024revisiting ; gu2023knowledge .
+Transformer-based large language models (LLMs) (Vaswani, 2017) have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks.
+This success is largely driven by their massive scale, often with billions of parameters.
+For instance, the OPT series (Zhang et al., 2022) includes models with up to 66 billion parameters, while the LLaMA family (Touvron et al., 2023a) features even larger models, such as LLaMA3-70B (Dubey et al., 2024).
+Although these models have substantially advanced the state of the art (SOTA) in NLP, their growing computational and memory demands present increasingly significant challenges for real-world deployment.
+The compression of LLMs can be broadly categorized into several approaches, including weight quantization (Lin et al., 2024; Frantar et al., 2023), low-rank factorization (Zhang et al., 2024; Yuan et al., 2023), network pruning (Sun et al., 2024; Frantar & Alistarh, 2023), and knowledge distillation (Zhong et al., 2024; Gu et al., 2024).
 Among these, binarization represents an extreme form of quantization, reducing model weights to 1 bit and drastically lowering memory consumption.
-Recent efforts primarily focus on post-training quantization (PTQ) huang2024billm ; li2024arb ; dong2024stbllm , which eliminates the need for backpropagation, thereby accelerating the binarization process and improving deployment efficiency.
-For example, BiLLM huang2024billm introduces a residual approximation strategy to enhance the performance of 1-bit LLMs, while ARB-LLM li2024arb adopts an alternating refinement approach to better align binarized weights with their full-precision counterparts.
-STBLLM dong2024stbllm further advances this line of work by compressing LLMs to sub-1-bit precision.
+Recent efforts primarily focus on post-training quantization (PTQ) (Huang et al., 2024; Li et al., 2025; Dong et al., 2025), which eliminates the need for backpropagation, thereby accelerating the binarization process and improving deployment efficiency.
+For example, BiLLM (Huang et al., 2024) introduces a residual approximation strategy to enhance the performance of 1-bit LLMs, while ARB-LLM (Li et al., 2025) adopts an alternating refinement approach to better align binarized weights with their full-precision counterparts.
+STBLLM (Dong et al., 2025) further advances this line of work by compressing LLMs to sub-1-bit precision.
 Despite the aforementioned advancements, binarized LLMs continue to exhibit redundancy, indicating opportunities for additional compression.
-Pruning lecun1989optimal is a promising technique for further reducing redundancy in binarized models.
-However, conventional structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation often leads to severe performance degradation in LLMs.
-On the other hand, unstructured pruning dong2024pruner suffers from poor hardware compatibility and inefficient storage.
-Semi-structured pruning frantar2023sparsegpt ; sun2023simple ; dong2024stbllm has emerged as an effective compromise, significantly reducing redundancy while maintaining a balance between accuracy and hardware efficiency.
-Nonetheless, directly combining binarization with semi-structured pruning often leads to noticeable performance degradation, making it hard to maintain model accuracy.
-Another challenge is how to select pruning elements effectively to improve pruning efficiency while preserving accuracy.
+Pruning (LeCun et al., 1989) is a promising technique for further reducing redundancy in binarized models.
+However, conventional structured pruning (Ma et al., 2023; Ashkboos et al., 2024; Xia et al., 2024; An et al., 2024) often leads to severe performance degradation in LLMs.
+On the other hand, unstructured pruning (Dong et al., 2024) suffers from poor hardware compatibility and inefficient storage.
+Semi-structured pruning (Frantar & Alistarh, 2023; Sun et al., 2024; Dong et al., 2025) has emerged as an effective compromise, significantly reducing redundancy while maintaining a balance between accuracy and hardware efficiency.
+Nonetheless, directly combining binarization with semi-structured pruning often leads to noticeable performance degradation, as the joint constraints on weight values and structure increase the difficulty of preserving model accuracy under extreme compression.
+This problem is further complicated by the need to select pruning elements carefully, since suboptimal choices may remove important weights and amplify quantization errors.
+Therefore, designing a unified framework that balances compression and accuracy retention remains challenging.
 To address these challenges, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLMs, which achieves substantial model compression while maintaining strong performance (see Figure 1).
@@
 This makes it a provably optimal strategy for pruning.
+Together, SPBO and CFS form the core of PBS2P, enabling provably efficient compression with minimal accuracy degradation.
 Extensive experiments demonstrate that PBS2P consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing binary PTQ methods on a wide range of evaluation benchmarks.
-As shown in Figure 1, on the WikiText-2 merity2016pointer dataset, PBS2P attains a perplexity of 6.20 on LLaMA-13B touvron2023llama1 with an average bit-width of just 0.8 bits, compared to 5.47 for the full-precision model.
+As shown in Figure 1, on the WikiText-2 (Merity et al., 2017) dataset, PBS2P attains a perplexity of 6.20 on LLaMA-13B (Touvron et al., 2023a) with an average bit-width of just 0.8 bits, compared to 5.47 for the full-precision model.
 These results further highlight the effectiveness of PBS2P in significantly narrowing the performance gap between binarized and full-precision models.
@@
 Current quantization methods for LLMs are mainly divided into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ).
-QAT liu2023llm ; chen2024db ; du2024bitdistiller integrates quantization during the training phase to enhance low-bit weight representations.
+QAT (Liu et al., 2024; Chen et al., 2024; Du et al., 2024) integrates quantization during the training phase to enhance low-bit weight representations.
 However, due to the enormous parameter number, retraining becomes excessively expensive and inefficient for LLMs.
 PTQ, as it directly applies quantization to the model weights without retraining, making it faster and less resource-demanding.
-Recent methods, like ZeroQuant yao2022zeroquant and BRECQ li2021brecq , improve quantization accuracy by incorporating custom quantization blocks and group labels.
-While GPTQ frantar2022gptq and QuIP chee2024quip use second-order error compensation to reduce quantization errors.
+Recent methods, like ZeroQuant Yao et al.
+(2022) and BRECQ (Li et al., 2021), improve quantization accuracy by incorporating custom quantization blocks and group labels.
+While GPTQ (Frantar et al., 2023) and QuIP (Chee et al., 2024) use second-order error compensation to reduce quantization errors.
 2.2 Network Binarization
 Binarization, as the most extreme form of quantization, reduces model parameters to a single bit (±1).
-Prominent methods, like Binary Weight Network (BWN) rastegari2016xnor , XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi , focus on binarizing the weights, with XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi also binarizing activations.
-In the context of LLM binarization, BitNet wang2023bitnet , OneBit (xu2024onebit, ), and BinaryMoS jo2024mixture adopt the QAT framework, while BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLMdong2024stbllm use PTQ combined with residual approximation.
-Our work focuses on the binary PTQ method, achieving substantial improvements over existing SOTA binary PTQ methods.
+Prominent methods, like Binary Weight Network (BWN) (Rastegari et al., 2016), XNOR-Net (Rastegari et al., 2016) and Bi-Real Net(Liu et al., 2018), focus on binarizing the weights, with XNOR-Net (Rastegari et al., 2016) and Bi-Real Net(Liu et al., 2018) also binarizing activations.
+In the context of LLM binarization, BitNet (Wang et al., 2023), OneBit (Xu et al., 2024), and BinaryMoS (Jo et al., 2024) adopt the QAT framework, while BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), and STBLLM(Dong et al., 2025) use PTQ combined with residual approximation.
+Focusing on the PTQ setting, our method yields significant enhancements over prior SOTA binary PTQ techniques.
 2.3 Model Pruning
-Pruning lecun1989optimal is a common technique for compressing neural networks by removing less important parameters, resulting in smaller and more efficient sparse models.
+Pruning (LeCun et al., 1989) is a common technique for compressing neural networks by removing less important parameters, resulting in smaller and more efficient sparse models.
 In LLMs, pruning methods are typically categorized as structured, unstructured, or semi-structured.
-Structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation removes entire components to enhance efficiency but often causes performance drops, requiring retraining.
-Unstructured pruning dong2024pruner eliminates individual weights based on importance, preserving performance even at high sparsity, though the irregular patterns are hardware-unfriendly.
-Semi-structured pruning offers a good trade-off, enforcing regular patterns like $N$:$M$ sparsity for hardware efficiency, as used in SparseGPT frantar2023sparsegpt , Wanda sun2023simple , and STBLLM dong2024stbllm .
-Our method adopts semi-structured $N$:$M$ sparsity to retain performance while ensuring hardware compatibility.
+Structured pruning (Ma et al., 2023; Ashkboos et al., 2024; Xia et al., 2024; An et al., 2024) removes entire components to enhance efficiency but often causes performance drops, requiring retraining.
+Unstructured pruning (Dong et al., 2024) eliminates individual weights based on importance, preserving performance even at high sparsity, though the irregular patterns are hardware-unfriendly.
+Semi-structured pruning offers a good trade-off, enforcing regular patterns like NN:$M$ sparsity for hardware efficiency, as used in SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), and STBLLM (Dong et al., 2025).
+We adopt NN:$M$ sparsity to balance performance and hardware efficiency.
 Figure 2: Overview of our PBS2P framework.
@@
 Binarization.
-First, let’s briefly review the process of standard binarization rastegari2016xnor .
-Binarization begins by performing a row-wise redistribution on the full-precision matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ , resulting in a matrix $\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}$ with a mean of 0:
-$\widetilde{\mathbf{W}}=\mathbf{W}-\mu,\quad\text{where}\,\mu=\frac{1}{m}\sum_{j=1}^{m}\mathbf{W}_{.j}\,.$ (1)
-The binarization objective is to minimize the reconstruction error: $\arg\min_{\alpha,\mathbf{B}}\|\widetilde{\mathbf{W}}-\alpha\mathbf{B}\|_{F}^{2}$, where $\alpha\in\mathbb{R}^{n}$ is a row-wise scaling factor and $\mathbf{B}\in\{+1,-1\}^{n\times m}$ is the binary matrix.
-The optimal solutions for $\alpha$ and $\mathbf{B}$ are given by $\alpha=\frac{1}{m}\sum_{j=1}^{m}|\widetilde{\mathbf{W}}_{.j}|$ and $\mathbf{B}=\operatorname{sign}(\widetilde{\mathbf{W}})$, respectively huang2024billm .
+We begin by revisiting the standard binarization process (Rastegari et al., 2016).
+The full-precision matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ is first row-wise normalized by subtracting the mean of each row, yielding a zero-centered matrix $\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}$:
+$\widetilde{\mathbf{W}}=\mathbf{W}-\mu,\quad\text{where}\quad\mu=\frac{1}{m}\sum_{j=1}^{m}\mathbf{W}_{.j}\,.$ (1)
+[…]
 N:M Sparsity.
-$N$:$M$ sparsity is a semi-structured pruning scheme that balances model compression and hardware efficiency.
-Unlike unstructured pruning, which removes arbitrary weights and results in irregular sparsity patterns that are hard to accelerate, $N$:$M$ sparsity imposes a structured constraint by retaining exactly $N$ nonzero elements out of every $M$ consecutive weights frantar2023sparsegpt ; sun2023simple ; dong2024stbllm .
+NN:$M$ sparsity is a semi-structured pruning scheme that balances model compression and hardware efficiency.
+Unlike unstructured pruning, which removes arbitrary weights and results in irregular sparsity patterns that are hard to accelerate, NN:$M$ sparsity imposes a structured constraint by retaining exactly $N$ nonzero elements out of every $M$ consecutive weights (Frantar & Alistarh, 2023; Sun et al., 2024; Dong et al., 2025).
 A typical configuration, such as 2:4 sparsity, keeps only two weights out of every four, significantly reducing parameter count while preserving computational efficiency.
-This regular pattern allows for efficient execution on NVIDIA Ampere GPUs nvidia2020 , taking advantage of specialized hardware support.
-In our work, we adopt the $N$:$M$ sparsity scheme as a semi-structured pruning strategy to further eliminate redundancy within the binarized model.
-This combination enables more aggressive compression while maintaining hardware compatibility and inference efficiency.
+This regular pattern allows for efficient execution on NVIDIA Ampere GPUs (Nvidia, 2020), taking advantage of specialized hardware support.
+In our work, we adopt the NN:$M$ sparsity scheme as a semi-structured pruning strategy to further eliminate redundancy within the binarized model.
+This enables stronger compression with hardware-friendly execution.
 3.2 Stepwise Semi-Structured Pruning with Binarization Optimization
-We define the total error $\mathcal{L}$ for binarization and semi-structured pruning as follows:
-$\mathcal{L}=||\mathbf{W}\mathbf{X}-(\mathbf{M}_{\text{p}}\odot\mathbf{\widehat{W}})\mathbf{X}||_{F}^{2},\quad\text{where}\,\widehat{\mathbf{W}}=\alpha\mathbf{B}+\mu\,.$ (2)
+To facilitate the joint optimization of binarization and semi-structured pruning, we formulate the total compression error $\mathcal{L}$ as the objective function to be minimized:
+[…]
 where $\mathbf{X}$ denotes the calibration data, $\mathbf{M}_{\text{p}}$ denotes the semi-structured pruning mask.
-Optimizing both the pruning mask $\mathbf{M}_{\text{p}}$ and the binarized parameters $\mathbf{\widehat{W}}$ simultaneously is an NP-hard problem blumensath2008iterative .
+Optimizing both the pruning mask $\mathbf{M}_{\text{p}}$ and the binarized parameters $\mathbf{\widehat{W}}$ simultaneously is an NP-hard problem (Blumensath & Davies, 2008).
 Therefore, we adopt a greedy-style approximation.
@@
 This step-by-step strategy reduces the optimization difficulty.
-It distributes the optimization burden across multiple pruning stages.
-The selection of the semi-strucutred pruning mask $\mathbf{M}_{\text{p}}$ will be elaborated in Section 3.3.
-Figure 3: Illustration of the $N$:$M$ sparsity pruning mask group, where the original pruning mask $\mathbf{M}_{\text{p}}$ is divided into $M$-$N$ sub masks.
+It distributes optimization across multiple pruning stages.
+Selection of the semi-structured mask $\mathbf{M}_{\text{p}}$ is detailed in Section 3.3.
+Figure 3: Illustration of the NN:$M$ sparsity pruning mask group, where the original pruning mask $\mathbf{M}_{\text{p}}$ is divided into $M$-NN sub masks.
 N: M Sparsity Pruning Mask Group.
-To enable progressive pruning, we decompose the original pruning mask $\mathbf{M}_{\text{p}}$ into a series of sub-masks, each representing a step of pruning.
-Under the $N$:$M$ sparsity constraint, exactly $N$ out of every $M$ elements are retained.
-Within each $M$-element group, the $M{-}N$ pruned elements are removed one by one across successive steps.
-Formally, the original pruning mask $\mathbf{M}_{\text{p}}$ can be expressed as the union of $M{-}N$ sub-masks:
-$\mathbf{M}_{\text{p}}=\bigcup_{i=1}^{M-N}\mathbf{M}_{\text{p}}^{i}.$ (3)
-Figure 3 illustrates this process. This decomposition supports progressive pruning, facilitating tighter integration with binarization in subsequent stages.
+To support progressive pruning, we decompose the original pruning mask $\mathbf{M}_{\text{p}}$ into a sequence of sub-masks, each corresponding to one intermediate pruning step.
+Under the NN:$M$ sparsity constraint, exactly $N$ out of every $M$ elements are retained.
+Within each $M$-element group, the MNM{-}N pruned elements are removed sequentially across successive stages.
+Formally, the full pruning mask can be expressed as the union of MNM{-}N sub-masks: $\mathbf{M}_{\text{p}}=\bigcup_{i=1}^{M-N}\mathbf{M}_{\text{p}}^{i}$.
+This decomposition enables fine-grained progressive pruning and allows for a tighter and more seamless integration with binarization in subsequent stages, as illustrated in Figure 3.
 Stepwise Pruning with Binarization Optimization (SPBO).
 After obtaining the group of pruning masks, we proceed with the stepwise pruning and binarization optimization process.
-We perform a total of $M{-}N$ pruning steps to achieve $N$:$M$ sparsity.
+We perform a total of MNM{-}N pruning steps to achieve NN:$M$ sparsity.
 After each step, we update the binarization parameters $\alpha$ and $\mu$ to better adapt to the current pruning state.
@@
 To minimize the error $\mathcal{L}$ in the current pruning state, these parameters must be re-estimated.
-Inspired by ARB-LLM li2024arb , a straightforward approach is to compute the partial derivatives of $\mathcal{L}$ with respect to $\alpha$ and $\mu$, and solve for the stationary points by setting them to zero.
+Inspired by Li et al.
+(2025), a straightforward approach is to compute the partial derivatives of $\mathcal{L}$ with respect to $\alpha$ and $\mu$, and solve for the stationary points by setting them to zero.
 Specifically, we first update $\mu$ by setting $\partial\mathcal{L}/\partial\mu=0$ during the $k$-th pruning step:
@@
 During the $k$-th pruning round, we can alternately update $\mu$ and $\alpha$ for $\tau$ steps to further reduce $\mathcal{L}$.
-We perform total $M$-$N$ steps of pruning until all pruning elements are removed.
+We perform total $M$-NN steps of pruning until all pruning elements are removed.
 The detailed derivation of optimizing $\mu$ and $\alpha$ can be found in the supplementary material.
 3.3 Coarse-to-Fine Search
-Selecting effective pruning elements is a key challenge in model compression.
+Selecting effective pruning elements is a key challenge in model pruning.
 To address this, we adopt a two-stage strategy.
@@
 Coarse-Stage: Pruning Ratio Allocation.
-To estimate the importance of each layer, we adopt a simple criterion that quantifies how much a layer influences the input-output relationship.
-Layers that induce greater changes between input and output are considered more important, whereas those with limited impact are deemed less important.
-Specifically, for a given weight layer $\mathcal{W}$, let $\mathcal{X}_{\mathcal{W}}$ and $\mathcal{Y}_{\mathcal{W}}$ denote the input and output, respectively.
-The importance of the layer is then evaluated as follows:
+To estimate the importance of each layer, inspired by Dumitru et al.
+(2024), we use cosine similarity to measure the alignment between its input and output representations.
+We adopt a straightforward criterion that quantifies how strongly a layer shapes the input–output relationship.
+Layers that induce larger changes are regarded as more important, whereas those with minimal effect are considered less important.
+Specifically, for a given weight layer $\mathcal{W}$, let $\mathcal{X}_{\mathcal{W}}$ and $\mathcal{Y}_{\mathcal{W}}$ denote its input and output, respectively.
+The importance of the layer is then computed as follows:
 $\text{Importance}(\mathcal{W})=\frac{1}{\text{similarity}(\mathcal{X_{\mathcal{W}},Y_{\mathcal{W}}})},\quad\text{where}\,\,\,\text{similarity}(\mathcal{X_{\mathcal{W}},Y_{\mathcal{W}}})=\frac{\mathcal{X}_{\mathcal{W}}\cdot\mathcal{Y}_{\mathcal{W}}}{\|\mathcal{X}_{\mathcal{W}}\|_{2}\|\mathcal{Y}_{\mathcal{W}}\|_{2}}.$ (6)
-Inspired by dumitru2024change , we employ cosine similarity to assess the alignment between input and output representations.
-Based on the computed importance, we rank all layers, with $k_{i}$ denoting the rank of the $i$-th layer.
-The pruning parameter $N_{i}$ for each layer is determined as follows:
+Based on the computed importance, we rank all layers, with $k_{i}$ denoting the rank of the $i$-th layer. The pruning parameter $N_{i}$ for each layer is determined as follows:
 $N_{i}=\left\lfloor N_{\text{high}}-\left(N_{\text{high}}-N_{\text{low}}\right)\cdot\frac{k_{i}-1}{L-1}+\frac{1}{2}\right\rfloor,$ (7)
@@
 We revisit the impact of pruning on model error.
-Inspired by OBS hassibi1993optimal , we derive the relationship between pruning and error increase under a second-order approximation, and present the corresponding formulation in Theorem 3.1.
+Inspired by Hassibi et al.
+(1993), we derive the relationship between pruning and error increase under a second-order approximation, and present the corresponding formulation in Theorem 3.1.
 The full proof is provided in the supplementary material.
 Since our objective is to minimize the increase in error caused by pruning, we directly adopt the following score to guide element selection: $s_{q}=\frac{w_{q}^{2}}{\left[\mathbf{H}^{-1}\right]_{qq}^{2}}$.
-This criterion is theoretically guaranteed to minimize the error increase introduced by pruning under a second-order approximation.Based on the given $N$:$M$ ratio and the corresponding $s_{q}$ values, we select the top $N$ elements with the largest $s_{i}$ values from every $M$ consecutive elements to retain.
-The other $M-N$ elements are selected for pruning.
+This criterion is theoretically guaranteed to minimize the error increase introduced by pruning under a second-order approximation.
+Based on the given NN:$M$ ratio and the corresponding $s_{q}$ values, we select the top $N$ elements with the largest $s_{i}$ values from every $M$ consecutive elements to retain.
+The other MNM-N elements are selected for pruning.
 Theorem 3.1.
-After pruning an element $w_{q}$, the pruning-induced error increase $\delta\mathcal{L}$ can be approximated by the following expression: $\delta\mathcal{L}=\frac{1}{2}\frac{w_{q}^{2}}{[\mathbf{H}^{-1}]_{qq}},$ (9) where $[\mathbf{H}^{-1}]_{qq}$ denotes the $q$-th diagonal entry of the inverse Hessian matrix.
+After pruning an element $w_{q}$, the pruning-induced error increase $\delta\mathcal{L}$ can be approximated by the following expression: $\delta\mathcal{L}=\frac{1}{2}\frac{w_{q}^{2}}{[\mathbf{H}^{-1}]_{qq}},$ (8) where $\mathbf{H}^{-1}]_{qq}$ denotes the $q$-th diagonal entry of the inverse Hessian matrix.
 Algorithm 1 Pseudocode of PBS2P
@@
 3:procedure PBS2P($\mathcal{M},\mathcal{X},\mathcal{TR},\tau$)
-4:     $R\leftarrow$ Coarse_Stage_Search$(\mathcal{M},\mathcal{X},\mathcal{TR})$triangleright Allocate layer-wise pruning ratios
+[…]
 5:  for each layer $l$ in model $\mathcal{M}$ do
-6:         $\mathcal{W}_{l}\leftarrow$ Extract_Weights$(\mathcal{M},l)$triangleright Obtain weights of layer $l$
-7:         $r_{l}\leftarrow R[l]$triangleright Get pruning ratio for layer $l$
-8:         $\mathbf{M}_{\text{p}}\leftarrow$ Fine_Stage_Search$(\mathcal{W}_{l},\mathcal{X},r_{l})$triangleright Generate pruning mask
-9:         $\mathcal{W}_{l}^{\prime}\leftarrow$ SPBO$(\mathcal{W}_{l},\mathcal{X},\mathbf{M}_{\text{p}},\tau)$triangleright Stepwise pruning and binarization optimization
-10:         $\mathcal{M}^{\prime}\leftarrow$ Update_Layer$(\mathcal{M},l,\mathcal{W}_{l}^{\prime})$triangleright Write updated weights back to model
+6:   $\mathcal{W}_{l}\leftarrow$ Extract_Weights(,l)($\mathcal{M}$,l) $\triangleright$ Obtain weights of layer $l$
+7:   $r_{l}\leftarrow R[l]$ $\triangleright$ Get pruning ratio for layer $l$
+[…]
+[…]
+[…]
 11:  end for
@@
 Average Bits.
-Following BiLLM huang2024billm and STBLLM dong2024stbllm , we similarly divide weights into salient and non-salient groups, consistent with their design.
+Following BiLLM (Huang et al., 2024) and STBLLM (Dong et al., 2025), we similarly divide weights into salient and non-salient groups, consistent with their design.
 For salient weights, we adopt a residual binarization strategy, allocating two bits to better preserve their expressiveness, as done in both prior works.
 For non-salient weights, we apply group-wise quantization.
-In line with STBLLM dong2024stbllm , we further divide the non-salient weights into three groups using two split points, maintaining the same quantization scheme.
+In line with STBLLM (Dong et al., 2025), we further divide the non-salient weights into three groups using two split points, maintaining the same quantization scheme.
 The weight parameters and additional hardware overhead are as follows:
-$\left\{\begin{array}[]{lr}N_{\text{param}}=\left[2\times r_{\text{salient}}+(1-r_{\text{salient}})\right]\times\frac{N}{M},\\ N_{\text{storing}}=2+\dfrac{1}{b_{\text{size}}},\\ \end{array}\right.$ (10)
-where $r_{\text{salient}}$ represents the proportion of salient weights, $\frac{N}{M}$ denotes the predefined avrage pruning ratio for the entire model, and $b_{\text{size}}$ indicates the block size used in OBC frantar2022optimal compensation, with 2 bits reserved to mark the division between salient and non-salient weights.
-Our parameter settings and $\frac{N}{M}$ configuration are identical to our main comparison method STBLLM dong2024stbllm .
-Specifically, we adopt the same settings of 4:8, 5:8, and 6:8 sparsity as used in STBLLM dong2024stbllm .
+[…]
+where $r_{\text{salient}}$ represents the proportion of salient weights, $\frac{N}{M}$ denotes the predefined average pruning ratio for the entire model, and $b_{\text{size}}$ indicates the block size used in OBC (Frantar & Alistarh, 2022) compensation, with 2 bits reserved to mark the division between salient and non-salient weights.
+Our parameter settings and $\frac{N}{M}$ configuration are identical to our main comparison method STBLLM (Dong et al., 2025).
+Specifically, we adopt the same settings of 4:8, 5:8, and 6:8 sparsity as used in STBLLM (Dong et al., 2025).
 These correspond to average bit-widths of 0.55, 0.70, and 0.80, respectively.
+This alignment ensures fair, direct comparison under consistent compression budgets.
 4 Experiments
 4.1 Settings
-All experiments are performed using PyTorch paszke2019pytorch and Huggingface paszke1912imperative on a single NVIDIA A800-80GB GPU.
-Following the work of frantar2022gptq , huang2024billm , and li2024arb , we use 128 samples from the C4 (raffel2020exploring, ) dataset for calibration.
+All experiments are performed using PyTorch (Paszke et al., 2019b) and Huggingface (Paszke et al., 2019a) on a single NVIDIA A800-80GB GPU.
+Following the work of (Frantar et al., 2023), (Huang et al., 2024), and (Li et al., 2025), we use 128 samples from the C4 (Raffel et al., 2020) dataset for calibration.
 Since PBS2P is an efficient PTQ framework, it eliminates fine-tuning, enabling completion through a single process combining binarization and pruning.
 Models and Datasets.
-We conduct extensive experiments on the LLaMA touvron2023llama1 , LLaMA-2 touvron2023llama2 , and LLaMA-3 dubey2024llama3 families and the OPT family (zhang2022opt, ).
-To evaluate the effectiveness of PBS2P, we measure the perplexity of LLM’s outputs on WikiText2 (merity2016pointer, ), PTB (marcus1994penn, ), and C4 (raffel2020exploring, ).
-Moreover, we also evaluate the accuracy for 7 zero-shot QA datasets: ARC-c (clark2018think, ), ARC-e (clark2018think, ), BoolQ (clark2019boolq, ), Hellaswag (zellers2019hellaswag, ), OBQA (mihaylov2018can, ), RTE (chakrabarty2021figurative, ), and Winogrande (sakaguchi2019adversarial, ).
+We conduct extensive experiments on the LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3 (Dubey et al., 2024) families and the OPT family (Zhang et al., 2022).
+To evaluate the effectiveness of PBS2P, we measure the perplexity of LLM outputs on WikiText2 (Merity et al., 2017), PTB (Marcus et al., 1994), and C4 (Raffel et al., 2020).
+Moreover, we also evaluate the accuracy on seven zero-shot QA datasets: ARC-c (Clark et al., 2018), ARC-e (Clark et al., 2018), BoolQ (Clark et al., 2019), Hellaswag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), RTE (Chakrabarty et al., 2021), and Winogrande (Sakaguchi et al., 2020).
+This benchmark setup validates our method on both language modeling and reasoning tasks.
 Baselines.
-We mainly compare our PBS2P with STBLLM dong2024stbllm , a structural binary PTQ framework designed for compressing LLMs to precisions lower than 1-bit.
-We compare the results of PBS2P with STBLLM under the same $N$:$M$ settings (e.g., 4:8, 5:8, 6:8).
-Previous low-bit methods like ARB-LLM li2024arb , BiLLM huang2024billm , GTPQ (frantar2022gptq, ) and vanilla RTN are also selected for comparison.
-Table 1: Perplexity comparison of RTN, GPTQ frantar2022gptq , BiLLM huang2024billm , ARB-LLM li2024arb , STBLLM dong2024stbllm , and PBS2P on the LLaMA families.
-The evaluation results demonstrate the perplexity performance on the Wikitext2 (merity2016pointer, ) dataset across various model sizes.
+We mainly compare our PBS2P with STBLLM (Dong et al., 2025), a structural binary PTQ framework designed for compressing LLMs to precisions lower than 1-bit.
+We compare the results of PBS2P with STBLLM under the same NN:$M$ sparsity settings (e.g., 4:8, 5:8, 6:8).
+Previous low-bit methods like ARB-LLM (Li et al., 2025), BiLLM (Huang et al., 2024), GTPQ (Frantar et al., 2023), and vanilla RTN are also selected as baselines for comparison.
+Table 1: Perplexity comparison of RTN, GPTQ (Frantar et al., 2023), BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), STBLLM (Dong et al., 2025), and PBS2P on the LLaMA families.
+The evaluation results demonstrate the perplexity performance on the Wikitext2 (Merity et al., 2017) dataset across various model sizes.
 Settings LLaMA-1 LLaMA-2 LLaMA-3 Method #Block W-Bits 7B 13B 30B 65B 7B 13B 8B FP16 - 16 5.68 5.09 4.10 3.53 5.47 4.88 6.14 RTN - 3 25.54 11.40 14.89 10.59 542.80 10.68 2194.98 GPTQ 128 3 8.63 5.67 4.87 4.17 6.44 5.46 18.68 RTN - 2 106767.34 57409.93 26704.36 19832.87 17788.94 51145.61 1335816.13 GPTQ 128 2 129.20 20.46 15.29 8.66 52.22 23.63 1480.43 RTN - 1 168388.00 1412020.25 14681.76 65253.24 157058.34 47902.32 1353698.38 GPTQ 128 1 164471.78 131505.41 10339.15 20986.16 59758.69 22926.54 1121260.50 BiLLM 128 1.11 49.79 14.58 9.90 8.37 32.31 21.35 55.80 ARB-LLM 128 1.11 14.03 10.18 7.75 6.56 16.44 11.85 27.42 STBLLM 128 0.80 15.03 9.66 7.56 6.43 13.06 11.67 33.44 STBLLM 128 0.70 19.48 11.33 9.19 7.91 18.74 13.26 49.12 STBLLM 128 0.55 31.72 17.22 13.43 11.07 27.93 20.57 253.76 PBS2P 128 0.80 7.36 6.20 5.21 4.60 7.17 6.27 10.75 PBS2P 128 0.70 8.09 6.85 5.78 5.09 8.00 6.89 12.29 PBS2P 128 0.55 10.78 9.24 7.19 6.39 10.64 8.68 17.45
 4.2 Main Results
-We perform a comprehensive comparison of different LLM families (like LLaMA touvron2023llama1 ; dubey2024llama3 ; touvron2023llama2 and OPT zhang2022opt ) with various model sizes.
-To keep fairness, We follow STBLLM dong2024stbllm to report the average bit-width of all methods, where our methods have the same bit-width as STBLLM.
+We perform a comprehensive comparison of different LLM families (like LLaMA (Touvron et al., 2023a; Dubey et al., 2024; Touvron et al., 2023b) and OPT (Zhang et al., 2022)) with various model sizes.
+To keep fairness, we follow STBLLM (Dong et al., 2025) to report the average bit-width of all methods, where our methods have the same bit-width as STBLLM.
 Results on LLaMA Family.
@@
 Figure 4 shows the performance of the LLaMA-7B, LLaMA-13B, and LLaMA-2-7B models.
-PBS2P continues to achieve a leading edge in performance while operating at a relatively lower bit-width compared to other methods.
+PBS2P continues to achieve a leading edge in performance while consistently operating at a relatively lower bit-width compared to other methods across diverse evaluation settings.
 Results on OPT Family.
 We extend our experiments to the OPT family (1.3B to 30B) under sub-1-bit PTQ settings, similar to the setup for LLaMA family.
-As shown in Table 2, PBS2P continues to outperform STBLLM across most of the models and $N$:$M$ structured binarization configurations.
+As shown in Table 2, PBS2P continues to outperform STBLLM across most of the models and NN:$M$ structured binarization configurations.
 More results are provided in the supplementary material.
@@
 Models Method W-Bits Winogrande OBQA Hellaswag Boolq ARC-e ARC-c RTE Average LLaMA-1-13B FP16 16 72.69 33.20 59.91 77.89 77.40 46.42 70.40 63.80 STBLLM 0.80 65.98 36.20 63.67 65.38 68.86 34.04 56.68 55.83 STBLLM 0.55 63.06 34.80 52.65 62.48 56.90 28.33 52.71 50.13 PBS2P 0.80 72.77 31.00 54.80 74.71 74.37 42.32 68.23 59.74 PBS2P 0.55 69.30 26.80 46.83 71.56 65.70 32.68 55.96 52.69 LLaMA-2-13B FP16 16 72.22 35.20 60.03 80.55 79.42 48.38 65.34 65.00 STBLLM 0.80 63.93 37.00 57.76 71.53 60.56 31.99 54.15 53.85 STBLLM 0.55 55.88 29.40 44.03 64.31 48.86 26.54 52.71 45.96 PBS2P 0.80 72.45 31.00 54.43 80.61 74.28 42.75 59.21 59.24 PBS2P 0.55 69.85 27.00 47.75 75.50 69.19 35.58 62.82 55.38 LLaMA-1-30B FP16 16 75.77 36.00 63.37 82.69 80.30 52.90 67.15 67.40 STBLLM 0.80 71.59 41.00 69.85 77.37 71.55 41.30 48.01 60.10 STBLLM 0.55 64.01 34.60 56.46 63.06 60.86 31.48 51.99 51.78 PBS2P 0.80 75.93 35.00 59.45 82.14 79.29 47.95 63.18 63.28 PBS2P 0.55 72.14 31.20 53.00 79.76 73.99 41.13 69.31 60.08
-Figure 4: BiLLM huang2024billm , ARB-LLM li2024arb , and PBS2P performed on the PTB (marcus1994penn, )and c4 raffel2020exploring datasets, mainly on LLaMA-7B, LLaMA-13B, and LLaMA2-7B, and we found that PBS2P at 0.8 bits outperforms other methods.
+Figure 4: BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), and PBS2P performed on the PTB (Marcus et al., 1994)and c4 (Raffel et al., 2020) datasets, mainly on LLaMA-7B, LLaMA-13B, and LLaMA2-7B, and we found that PBS2P at 0.8 bits outperforms other methods.
 4.4 Ablation Study
@@
 With the C4 dataset, reducing the sample size to 32 or 64 degrades performance, while increasing to 256 yields minimal additional benefit over 128.
-These results confirm that PBS2P remains robust with just 128 calibration samples.
-Table 4: Ablation study on LLaMA-7B, where all PBS2P is applied an $N$:$M$ sparsity of 4:8. Results are measured by perplexity on the Wikitext2 and C4 datasets. Our results are highlighted in bold.
-SPBO Strategy Wikitext2$\downarrow$ C4$\downarrow$
+PBS2P remains robust with only 128 calibration samples.
+Table 4: Ablation study on LLaMA-7B, where all PBS2P is applied an NN:$M$ sparsity of 4:8. Results are measured by perplexity on the Wikitext2 and C4 datasets. Our results are highlighted in bold.
+[…]
 14.43 16.76
@@
 Metric Random Magnitude Wanda SI Ours
-Wikitext2$\downarrow$ 7,779.28 38.90 10.89 196.61 10.78
-C4$\downarrow$ 6,797.09 24.47 13.16 148.85 13.06
+[…]
+[…]
 (b) Ablation for Metric in Fine-Stage Search
-Coarse-Stage Search Metric Wikitext2$\downarrow$ C4$\downarrow$
+[…]
 11.55 13.94
@@
 (c) Ablation for Metric in Coarse-Stage Search
-Prune Type Hardware-friendly Wikitext2$\downarrow$ C4$\downarrow$
+[…]
 Structured 621.16 400.50
@@
 Group Size 32 64 128 256 512
-Wikitext2$\downarrow$ 9.74 10.24 10.78 11.20 11.91
-C4$\downarrow$ 11.75 12.50 13.06 13.68 14.51
+[…]
+[…]
 (e) Ablation for Group Size
 Calibration Data Size 32 64 128 256
-Wikitext2$\downarrow$ 11.62 11.20 10.78 10.77
-C4$\downarrow$ 13.58 13.44 13.06 13.12
+[…]
+[…]
 (f) Ablation for Calibration Data Size
@@
 It introduces coarse-to-fine search and stepwise pruning with binarization optimization, which moderately increases the overall computation time.
-As shown in Table 5, BiLLM huang2024billm completes quantization on LLaMA-7B in 45 minutes, while ARB-LLM li2024arb takes 76 minutes.
+As shown in Table 5, BiLLM (Huang et al., 2024) completes quantization on LLaMA-7B in 45 minutes, while ARB-LLM (Li et al., 2025) takes 76 minutes.
 Table 5: Time comparison of different PTQ methods on LLaMA-7B.
@@
 With this design, PBS2P achieves superior accuracy and highly efficient, practical deployment.
-5 Discussion
-Limitations.
-While PBS2P demonstrates strong performance in both accuracy and model compression, it still faces several limitations.
-First, we have not yet implemented a custom 1-bit kernel that supports the $N$:$M$ sparsity pattern, which limits our ability to accurately evaluate the end-to-end speedup on real hardware platforms.
-Second, similar to prior binarization methods such as BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLM dong2024stbllm , our approach introduces additional bits to differentiate between salient and non-salient weights.
-Although this selective binarization strategy helps preserve model performance, it also incurs extra storage overhead, which partially offsets the compression and efficiency gains brought by binarization.
-Societal impacts.
-Our proposed PBS2P method enables the efficient deployment of large language models on resource-constrained devices, promoting broader access to AI technologies in real-world applications such as mobile assistants, edge computing, and low-power systems.
-At present, we foresee no negative societal impacts arising from this work.
+Figure 5: Latency comparison (ms) of matrix multiplications across different projection shapes.
+4.6 Inference Efficiency
+We evaluate inference efficiency using custom matrix multiplication kernels for binarized and semi-structured sparse weights, built on top of the open-source BitBLAS library111 https://github.com/microsoft/BitBLAS.
+As shown in the figure 5, PBS²P (4:8) consistently achieves the lowest latency across typical LLM projection shapes, with activation sequence length fixed at 2048.
+Compared to FP16 and other 1-bit baselines like BiLLM and ARB-LLM, PBS²P delivers significant speedups, especially on wider matrices such as 4096×11008 and 11008×4096.
 5 Conclusion
@@
 Extensive experiments demonstrate that PBS2P outperforms SOTA binary PTQ methods, delivering superior accuracy across various LLM families and evaluation metrics.
-Our method provides an effective solution for deploying LLMs on resource-constrained devices while maintaining high performance, and offers a novel perspective on combining compression techniques for extreme LLM compression.
+Our method enables efficient LLM deployment on resource-limited devices while preserving performance and unifying compression strategies for extreme model reduction.